OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, 1) and (–2, 5)? A. 4/5 B. -5/4 C.-4/5 D. 5/4

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }=\frac{ 1-5 }{ 3+2 }=?\]

OpenStudy (anonymous):

-4/5 :) thanks

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) \(\Huge\color{blue}{☞}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/5655f2d1e4b0af360a7cfaa3 \(\Huge\color{blue}{☞}\) Consider reading http://openstudy.com/code-of-conduct and http://openstudy.com/terms-and-conditions to learn rules.